# StackTrans: From Large Language Model to Large Pushdown Automata Model

### v1

@@ 1 Introduction
In the era of large language models (LLMs) [GPT-4, 2023], the Transformer architecture has emerged as the nearly universal backbone, achieving remarkable success across various domains and beyond [Bi et al., 2024; Bai et al., 2023; Zhang et al., 2024a]. However, recent empirical studies [Delétang et al., 2022; Hahn, 2020] have demonstrated that Transformers struggle with tasks involving Chomsky hierarchies [Chomsky, 1956], such as regular expressions (REs) and deterministic context-free grammars (DCFs). For example, in an RE matching task, each training example consists of a string within a specified length range, paired with a matching-or-not label. While Transformers can perform well within the length boundaries of the training set, they often fail to maintain consistent performance when evaluated on input strings that are longer or shorter than those in the training data. One theoretical explanation is that standard Transformers lack inductive biases [Mitchell, 1980; Battaglia et al., 1806; Sartran et al., 2022]. Without inherent assumptions to guide learning, Transformers struggle to effectively capture the underlying grammar of the Chomsky hierarchy in the training set. Consequently, they cannot generalize beyond the training data and perform poorly on inputs of lengths that differ from those seen during training. On the other hand, natural languages are generally considered to belong to a class of languages that extends beyond DCF [Gazdar and Pullum, 1982; Chomsky, 1956; Shieber, 1985]. This inherent limitation of Transformers may possibly hinder their application in real-world natural language modeling. Such deficiencies may also have the risk to block LLMs to achieve more advanced forms of intelligence.
@@
In the realm of formal language theory and computational linguistics [Chomsky, 1956; Savage and Computation, 1998; Sipser, 1996], it is widely recognized that automata augmented with stacks correspond to different levels of the Chomsky hierarchy grammars  [^1] The Chomsky grammars of type 3 (REs) and type 2 (DCFs) can be recognized by pushdown automata (automata equipped with a stack) [Chomsky, 1956; Chomsky and Schützenberger, 1959]. With proper enhancement, pushdown automata can even resolve type 1 (context-sensitive, CS) and type 0 (unrestricted) grammars – 2-stack automata (automata equipped with two independent stacks) are equivalent to Turing machines, and can recognize type 0 grammars [Yau, 1969].. Given the success of stack-equipped automata in handling rather complex grammars, it is a natural progression to incorporate the data structure of stack into the Transformer architecture. Recently, researchers have proposed the stacked attention mechanism [DuSell and Chiang, 2024; Li et al., 2024] and have examined its viability upon multiple relatively small benchmarks.
@@
Drawing inspiration from pushdown automata and prior research, we introduce hidden state stacks into the Transformer architecture, proposing a novel method, StackTrans. Unlike the stacked attention mechanisms [DuSell and Chiang, 2024; Li et al., 2024], which replace the standard attention, StackTrans incorporates differentiable stacks between the Transformer layers, meanwhile preserving the integrity of the Transformer layers (Figure LABEL:fig:method:a). This integration allows us to embed the assumptions of the Chomsky hierarchy into the model, enabling StackTrans to inherently model and learn REs and DCFs. Moreover, this design maintains compatibility with existing frameworks like flash-attention [Dao et al., 2022; Dao, 2024], enabling seamless integration into efficient LLM training pipelines. Specifically, the stack stores hidden states generated by the preceding layer and updates through operations such as push and pop at each decoding time step. To enable end-to-end training of StackTrans, we design soft stack operations, thereby making the hidden state stack differentiable. StackTrans also adopts a multi-head stack to improve its representation capability. Additionally, we find that the standard stack reading operation (which only returns the top element in the stack) may cause unstable training. Therefore, we propose the global reading operation through a learnable query-over-stack attention, stabilizing the training process and enriching the expressiveness of StackTrans.
@@
We conduct comprehensive experiments on multiple benchmarks spanning both formal languages [Delétang et al., 2022] and natural languages [Groeneveld et al., 2024]. Specifically, in RE and DCF tasks, StackTrans outperforms the standard Transformer by at least 30%, achieving nearly 100% test accuracy in most scenarios, as shown in Figure LABEL:fig:method:b. Surprisingly, during natural language evaluation, StackTrans also demonstrates substantial improvements on tasks such as common sense reasoning and question answering. Furthermore, we have successfully scaled StackTrans up from 360M to 7B parameters. In particular, our open-sourced StackTrans-360M, which is pretrained on a corpus of approximately 1T tokens, performs better than or comparably to state-of-the-art LLMs with 2–3$\times$ more parameters, as shown in Figure LABEL:fig:method:c.
@@ 2 Background
The foundational success of LLMs can be attributed to the development of the Transformer architecture [Vaswani et al., 2017] and its numerous variations, which serve as the backbone for LLMs [Radford et al., 2018; GPT-4, 2023; GPT-4o, 2024]. Despite their widespread adoption, the Transformer architecture has inherent expressivity limitations [Hahn, 2020]. Recent studies have shown that standard Transformers struggle with recursive and hierarchical patterns across both synthetic and real-world tasks [Joulin and Mikolov, 2015; Grefenstette et al., 2015; Sartran et al., 2022], and need to equip neural networks with external data structures. These limitations pose a potential risk to the natural language modeling capabilities of Transformers, as suggested by discussions surrounding the classification of natural languages within the Chomsky hierarchy [Gazdar and Pullum, 1982; Chomsky, 1956; Shieber, 1985]. For more detailed related work, please refer to §A. Based on the same principle, StackTrans introduces differentiable hidden state stacks in a modular and scalable manner. Without altering the attention mechanism, StackTrans integrates stack operations into layer-wise hidden state updates. This design maintains compatibility with existing frameworks like flash-attention [Dao et al., 2022; Dao, 2024] and supports architectures ranging from 0.36B to 7B parameters. By learning stack operations explicitly, StackTrans is capable of addressing broader linguistic and algorithmic challenges.
@@ 3 Method
As introduced previously, the standard Transformer architecture [Vaswani et al., 2017] struggles to learn the Chomsky hierarchy due to its lack of inductive biases. To address this limitation, we propose StackTrans, which incorporates hidden state stacks into the Transformer architecture. These stacks introduce the assumptions of the Chomsky hierarchy, enhancing the model’s ability to capture hierarchical structures. In StackTrans, the hidden state stacks augment the information flow by routing token-level hidden states through learnable stack operations, such as stack updating and reading. Specifically, the stacks are integrated between standard Transformer layers, where they store hidden states and perform updates and readings via soft operations (see §3.1 for details). To further improve the expressiveness of StackTrans, we introduce a multi-head stack mechanism (please refer to §3.2). This design enhances the ability of StackTrans to capture diverse patterns with low-rank representations. Finally, to ensure robust training and avoid issues such as stack operation collapsing, we introduce stack regularization techniques. We also implement stack truncation to facilitate batching and parallel training (§3.3). These innovations and designs collectively enhance the effectiveness and stability of StackTrans in learning Chomsky hierarchy grammars.
@@ 3.1 Hidden State Stack
In computational linguistics, it is well established that the Chomsky hierarchy grammars can be resolved by different classes of automata, with pushdown automata being the minimal computational model for DCFs [Chomsky and Schützenberger, 1959]. Drawing inspiration from this, we incorporate stacks into the Transformer architecture to address its limitations in learning the Chomsky hierarchy. Recall that a standard stack is a last-in-first-out storage structure that allows the top element to be read or updated (i.e., operations like peek, push, and pop). Given a hidden state sequence $h_{1},\cdots,h_{l}$ (for now, we do not consider any computational dependencies among $h_{t}$’s), at each time-step $t$, StackTrans is designed to either push the current hidden state $h_{t}$ into the stack or pop the top element from the stack, and then peek the current top element (at this stage, we ignore how these operations are determined or executed by the model). This procedure results in a new stack-operated sequence, which is a permutation of $h_{1},\cdots,h_{l}$. Ideally, if the stack operations are correct, it is highly plausible that the model can effectively learn the target Chomsky hierarchy grammar.
Since the hidden state stack will be incorporated into StackTrans, it must be differentiable to enable end-to-end training. However, standard stack operations such as push and pop are discrete, which disrupts gradient back-propagation and hinders the training process. To address this challenge, we introduce a soft operation mechanism, following earlier explorations [Grefenstette et al., 2015; Joulin and Mikolov, 2015]. In this mechanism, the results of the stack operations are continuously interpolated based on some trainable parameters. This design not only makes the operations differentiable but also allows the model to learn the stack operations through these trainable parameters.
@@ 3.2 Multi-Head Low-Rank Stack
In the Transformer architecture, the multi-head attention mechanism [Vaswani et al., 2017] processes multiple attention patterns in parallel across different representation subspaces. The design enables the model to capture diverse relationships within the input sequence. Following a similar design philosophy, we propose the multi-head low-rank stack. Specifically, we down-project the hidden state $h_{t}\in\mathbb{R}^{d}$ and split it into subspaces ($\mathbb{R}^{d_{s}}$) as: $\left[h^{(1)}_{t},h^{(2)}_{t},\ldots,h^{(H)}_{t}\right]=\text{Reshape}(W_{\mathrm{down}}\cdot h_{t})$, where $H$ denotes the number of stack heads and $W_{\text{down}}\in\mathbb{R}^{(H\cdot d_{s})\times d}$ is the down-projection matrix. Each head corresponds to an independent stack as introduced in §3.1.
@@ 3.3 Key Implementation & Training Know-Hows
Ideally, the stack is supposed to process hidden states according to their temporal or layer dependencies, i.e., it should prioritize hidden states generated from earlier tokens or shallower layers. One feasible sequence fed into the hidden state stack would be $\left[h_{{t_{0}},0},h_{{t_{0}},1},\cdots,h_{{t_{0}},L},h_{{t_{1}},0},\cdots\right]$, where $h_{t,i}\in\mathbb{R}^{d}$ denotes the hidden state of the $i$-th Transformer layer at token $t$, and $L$ represents the total number of layers. Although such a behavioral pattern is clearly beneficial for learning stack operations, the temporal dependencies conflict with the parallel training of the Transformer layers. To facilitate training parallelism, we implement StackTrans by breaking these temporal dependencies. Specifically, StackTrans learns stack operations based on the hidden state sequence at token $t_{i}$ from layer 00 to $L$ ($\left[h_{{t_{i}},0},\cdots,h_{{t_{i}},L}\right]$), allowing all tokens to be trained in parallel. We provide detailed discussion in §B.3.
@@
    Results listed here are reported by DuSell and Chiang [2024].
@@ 4 Evaluation against Formal Languages
In this section, we evaluate StackTrans on three groups formal language modeling tasks aligned with the Chomsky hierarchy [Delétang et al., 2022]. These tasks assess a model’s ability to learn underlying compositional rules of formal languages and generalize to input lengths beyond those seen during training. Please refer to §D for details. Following prior work [Delétang et al., 2022], we implement StackTrans with relatively limited parameters. Concretely, we use five Transformer layers with $d=64$. $H$ is set to 4 and $d_{s}$ is set to 8. We compare StackTrans to some representative baselines, including the standard Transformer [Vaswani et al., 2017], LSTM [Graves and Graves, 2012], StackRNN [Joulin and Mikolov, 2015]) and StackAttn [Li et al., 2024], maintaining identical experimental settings for all models.
@@ 5 Evaluation against General Natural Languages
From the perspective of computational linguistics, natural languages are generally considered to belong to a class of languages that includes DCF [Gazdar and Pullum, 1982; Chomsky, 1956; Shieber, 1985]. Therefore, to thoroughly assess StackTrans, we conduct further evaluations on general language modeling tasks. ❶ We examine the scalability of StackTrans through scaling law studies, analyzing the effect caused by model size and training tokens. ❷ A StackTrans with 360 million parameters is pretrained with nearly 980 billion tokens. We evaluate its performance on a variety of standard benchmarks, comparing it to models with similar or larger parameter scales. ❸ We provide additional empirical observations through ablation studies and deeper analysis (please refer to §6).
@@
We follow the OLMo framework [allenai, 2024] to pretrain StackTrans. Our corpora comes from Dolma [Soldaini et al., 2024] and Smoll [Allal et al., 2025], which contain high-quality natural language, math, and Python code examples with diverse domains. We carry out data filtration, ultimately obtaining approximately 980 billion tokens. To scale up model parameters, we adapt the Dolma v1.6-sample configuration in OLMo, using roughly 80 billion tokens for each variant model training. For scaling up training tokens, we train StackTrans with 360M parameters on a sampled subset of 200 billion tokens from pretraining corpora.
@@ 6 Discussion
We evaluate all the variants introduced above on the V2 and V3 validation sets [Zhu et al., 2024]. Experimental details are provided in §G. The ablation results are presented in Table 3. The QueueTrans variant exhibits notably higher perplexity and lower overall accuracy compared to StackTrans, particularly on tasks involving hierarchical or recursive patterns. This outcome is consistent with our expectation that queue operations are inherently less effective at modeling nested dependencies and grammars. Similarly, the push-only variant performs poorly, with both training and validation losses significantly deteriorating. The absence of pop operations impairs its ability to dynamically manage and retrieve stored information, thereby reducing its overall effectiveness. Both single-head and full-dimensional variants are consistently outperformed by StackTrans. The multi-head mechanism enhances flexibility by enabling parallel decomposition of stack streams, while the low-rank mechanism reduces computational costs without compromising much modeling capacity.
@@
In addition to the configuration of the standard Transformer layers, StackTrans has three key hyperparameters – the number of stack heads $H$, the dimension of each stack head $d_{s}$, and the stack size $S$. We perform grid search over reasonable ranges for these hyperparameters, training StackTrans with 20 billion tokens. The curves of training loss and validation loss (evaluated on V2 [Zhu et al., 2024]) are plotted in Figure LABEL:fig:stackablation. It is clear that $H$ is crucial for parallelism, but Figure LABEL:fig:ablation:head indicates that performance plateaus once $H$ surpasses a certain threshold. From Figure LABEL:fig:ablation:dim, we observe that setting $d_{s}$ within the range from 16 to 64 balances the computational cost and the model’s expressiveness effectively. Similarly, Figure LABEL:fig:ablation:slots shows that increasing $S$ from 24 to 32 has nearly no impact on performance. Given that a larger $S$ leads to higher storage overhead, we make a trade-off and ultimately set $S$ to 24 during our evaluation.

### v2

@@ 1 Introduction
In the era of large language models (LLMs) (GPT-4, 2023), the Transformer architecture has emerged as the nearly universal backbone, achieving remarkable success across various domains and beyond (Bi et al., 2024; Bai et al., 2023; Zhang et al., 2024a). However, recent empirical studies (Delétang et al., 2022; Hahn, 2020) have demonstrated that Transformers struggle with tasks involving Chomsky hierarchies (Chomsky, 1956), such as regular expressions (REs) and deterministic context-free grammars (DCFs). For example, in an RE matching task, each training example consists of a string within a specified length range, paired with a matching-or-not label. While Transformers can perform well within the length boundaries of the training set, they often fail to maintain consistent performance when evaluated on input strings that are longer or shorter than those in the training data. One theoretical explanation is that standard Transformers lack inductive biases (Mitchell, 1980; Battaglia et al., 1806; Sartran et al., 2022). Without inherent assumptions to guide learning, Transformers struggle to effectively capture the underlying grammar of the Chomsky hierarchy in the training set. Consequently, they cannot generalize beyond the training data and perform poorly on inputs of lengths that differ from those seen during training. On the other hand, natural languages are generally considered to belong to a class of languages that extends beyond DCF (Gazdar and Pullum, 1982; Chomsky, 1956; Shieber, 1985). This inherent limitation of Transformers may possibly hinder their application in real-world natural language modeling. Such deficiencies may also have the risk to block LLMs to achieve more advanced forms of intelligence.
@@
In the realm of formal language theory and computational linguistics (Chomsky, 1956; Savage and Computation, 1998; Sipser, 1996), it is widely recognized that automata augmented with stacks correspond to different levels of the Chomsky hierarchy grammars  [^1] The Chomsky grammars of type 3 (REs) and type 2 (DCFs) can be recognized by pushdown automata (automata equipped with a stack) (Chomsky, 1956; Chomsky and Schützenberger, 1959). With proper enhancement, pushdown automata can even resolve type 1 (context-sensitive, CS) and type 0 (unrestricted) grammars – 2-stack automata (automata equipped with two independent stacks) are equivalent to Turing machines, and can recognize type 0 grammars (Yau, 1969).. Given the success of stack-equipped automata in handling rather complex grammars, it is a natural progression to incorporate the data structure of stack into the Transformer architecture. Recently, researchers have proposed the stacked attention mechanism (DuSell and Chiang, 2024; Li et al., 2024) and have examined its viability upon multiple relatively small benchmarks.
@@
Drawing inspiration from pushdown automata and prior research, we introduce hidden state stacks into the Transformer architecture, proposing a novel method, StackTrans. Unlike the stacked attention mechanisms (DuSell and Chiang, 2024; Li et al., 2024), which replace the standard attention, StackTrans incorporates differentiable stacks between the Transformer layers, meanwhile preserving the integrity of the Transformer layers (Figure LABEL:fig:method:a). This integration allows us to embed the assumptions of the Chomsky hierarchy into the model, enabling StackTrans to inherently model and learn REs and DCFs. Moreover, this design maintains compatibility with existing frameworks like flash-attention (Dao et al., 2022; Dao, 2024), enabling seamless integration into efficient LLM training pipelines. Specifically, the stack stores hidden states generated by the preceding layer and updates through operations such as push and pop at each decoding time step. To enable end-to-end training of StackTrans, we design soft stack operations, thereby making the hidden state stack differentiable. StackTrans also adopts a multi-head stack to improve its representation capability. Additionally, we find that the standard stack reading operation (which only returns the top element in the stack) may cause unstable training. Therefore, we propose the global reading operation through a learnable query-over-stack attention, stabilizing the training process and enriching the expressiveness of StackTrans.
@@
We conduct comprehensive experiments on multiple benchmarks spanning both formal languages (Delétang et al., 2022) and natural languages (Groeneveld et al., 2024). Specifically, in RE and DCF tasks, StackTrans outperforms the standard Transformer by at least 30%, achieving nearly 100% test accuracy in most scenarios, as shown in Figure LABEL:fig:method:b. Surprisingly, during natural language evaluation, StackTrans also demonstrates substantial improvements on tasks such as common sense reasoning and question answering. Furthermore, we have successfully scaled StackTrans up from 360M to 7B parameters. In particular, our open-sourced StackTrans-360M, which is pretrained on a corpus of approximately 1T tokens, performs better than or comparably to state-of-the-art LLMs with 2–3$\times$ more parameters, as shown in Figure LABEL:fig:method:c.
@@ 2 Background
The foundational success of LLMs can be attributed to the development of the Transformer architecture (Vaswani et al., 2017) and its numerous variations, which serve as the backbone for LLMs (Radford et al., 2018; GPT-4, 2023; GPT-4o, 2024). Despite their widespread adoption, the Transformer architecture has inherent expressivity limitations (Hahn, 2020). Recent studies have shown that standard Transformers struggle with recursive and hierarchical patterns across both synthetic and real-world tasks (Joulin and Mikolov, 2015; Grefenstette et al., 2015; Sartran et al., 2022), and need to equip neural networks with external data structures. These limitations pose a potential risk to the natural language modeling capabilities of Transformers, as suggested by discussions surrounding the classification of natural languages within the Chomsky hierarchy (Gazdar and Pullum, 1982; Chomsky, 1956; Shieber, 1985). For more detailed related work, please refer to §A. Based on the same principle, StackTrans introduces differentiable hidden state stacks in a modular and scalable manner. Without altering the attention mechanism, StackTrans integrates stack operations into layer-wise hidden state updates. This design maintains compatibility with existing frameworks like flash-attention (Dao et al., 2022; Dao, 2024) and supports architectures ranging from 0.36B to 7B parameters. By learning stack operations explicitly, StackTrans is capable of addressing broader linguistic and algorithmic challenges.
@@ 3 Method
As introduced previously, the standard Transformer architecture (Vaswani et al., 2017) struggles to learn the Chomsky hierarchy due to its lack of inductive biases. To address this limitation, we propose StackTrans, which incorporates hidden state stacks into the Transformer architecture. These stacks introduce the assumptions of the Chomsky hierarchy, enhancing the model’s ability to capture hierarchical structures. In StackTrans, the hidden state stacks augment the information flow by routing token-level hidden states through learnable stack operations, such as stack updating and reading. Specifically, the stacks are integrated between standard Transformer layers, where they store hidden states and perform updates and readings via soft operations (see §3.1 for details). To further improve the expressiveness of StackTrans, we introduce a multi-head stack mechanism (please refer to §3.2). This design enhances the ability of StackTrans to capture diverse patterns with low-rank representations. Finally, to ensure robust training and avoid issues such as stack operation collapsing, we introduce stack regularization techniques. We also implement stack truncation to facilitate batching and parallel training (§3.3). These innovations and designs collectively enhance the effectiveness and stability of StackTrans in learning Chomsky hierarchy grammars.
@@ 3.1 Hidden State Stack
In computational linguistics, it is well established that the Chomsky hierarchy grammars can be resolved by different classes of automata, with pushdown automata being the minimal computational model for DCFs (Chomsky and Schützenberger, 1959). Drawing inspiration from this, we incorporate stacks into the Transformer architecture to address its limitations in learning the Chomsky hierarchy. Recall that a standard stack is a last-in-first-out storage structure that allows the top element to be read or updated (i.e., operations like peek, push, and pop). Given a hidden state sequence $h_{1},\cdots,h_{l}$ (for now, we do not consider any computational dependencies among $h_{t}$’s), at each time-step $t$, StackTrans is designed to either push the current hidden state $h_{t}$ into the stack or pop the top element from the stack, and then peek the current top element (at this stage, we ignore how these operations are determined or executed by the model). This procedure results in a new stack-operated sequence, which is a permutation of $h_{1},\cdots,h_{l}$. Ideally, if the stack operations are correct, it is highly plausible that the model can effectively learn the target Chomsky hierarchy grammar.
Since the hidden state stack will be incorporated into StackTrans, it must be differentiable to enable end-to-end training. However, standard stack operations such as push and pop are discrete, which disrupts gradient back-propagation and hinders the training process. To address this challenge, we introduce a soft operation mechanism, following earlier explorations (Grefenstette et al., 2015; Joulin and Mikolov, 2015). In this mechanism, the results of the stack operations are continuously interpolated based on some trainable parameters. This design not only makes the operations differentiable but also allows the model to learn the stack operations through these trainable parameters.
@@ 3.2 Multi-Head Low-Rank Stack
In the Transformer architecture, the multi-head attention mechanism (Vaswani et al., 2017) processes multiple attention patterns in parallel across different representation subspaces. The design enables the model to capture diverse relationships within the input sequence. Following a similar design philosophy, we propose the multi-head low-rank stack. Specifically, we down-project the hidden state $h_{t}\in\mathbb{R}^{d}$ and split it into subspaces ($\mathbb{R}^{d_{s}}$) as: $\left[h^{(1)}_{t},h^{(2)}_{t},\ldots,h^{(H)}_{t}\right]=\text{Reshape}(W_{\mathrm{down}}\cdot h_{t})$, where $H$ denotes the number of stack heads and $W_{\text{down}}\in\mathbb{R}^{(H\cdot d_{s})\times d}$ is the down-projection matrix. Each head corresponds to an independent stack as introduced in §3.1.
@@ 3.3 Key Implementation & Training Know-Hows
Ideally, the stack is supposed to process hidden states according to their temporal or layer dependencies, i.e., it should prioritize hidden states generated from earlier tokens or shallower layers. One feasible sequence fed into the hidden state stack would be $\left[h_{{t_{0}},0},h_{{t_{0}},1},\cdots,h_{{t_{0}},L},h_{{t_{1}},0},\cdots\right]$, where $h_{t,i}\in\mathbb{R}^{d}$ denotes the hidden state of the $i$-th Transformer layer at token $t$, and $L$ represents the total number of layers. Although such a behavioral pattern is clearly beneficial for learning stack operations, the temporal dependencies conflict with the parallel training of the Transformer layers. To facilitate training parallelism, we implement StackTrans by breaking these temporal dependencies. Specifically, StackTrans learns stack operations based on the hidden state sequence at token $t_{i}$ from layer 0 to $L$ ($\left[h_{{t_{i}},0},\cdots,h_{{t_{i}},L}\right]$), allowing all tokens to be trained in parallel. We provide detailed discussion in §B.3.
@@
    Results listed here are reported by DuSell and Chiang (2024).
@@ 4 Evaluation against Formal Languages
In this section, we evaluate StackTrans on three groups formal language modeling tasks aligned with the Chomsky hierarchy (Delétang et al., 2022). These tasks assess a model’s ability to learn underlying compositional rules of formal languages and generalize to input lengths beyond those seen during training. Please refer to §D for details. Following prior work (Delétang et al., 2022), we implement StackTrans with relatively limited parameters. Concretely, we use five Transformer layers with $d=64$. $H$ is set to 4 and $d_{s}$ is set to 8. We compare StackTrans to some representative baselines, including the standard Transformer (Vaswani et al., 2017), LSTM (Graves and Graves, 2012), StackRNN (Joulin and Mikolov, 2015)) and StackAttn (Li et al., 2024), maintaining identical experimental settings for all models.
@@ 5 Evaluation against General Natural Languages
From the perspective of computational linguistics, natural languages are generally considered to belong to a class of languages that includes DCF (Gazdar and Pullum, 1982; Chomsky, 1956; Shieber, 1985). Therefore, to thoroughly assess StackTrans, we conduct further evaluations on general language modeling tasks. ❶ We examine the scalability of StackTrans through scaling law studies, analyzing the effect caused by model size and training tokens. ❷ A StackTrans with 360 million parameters is pretrained with nearly 980 billion tokens. We evaluate its performance on a variety of standard benchmarks, comparing it to models with similar or larger parameter scales. ❸ We provide additional empirical observations through ablation studies and deeper analysis (please refer to §6).
@@
We follow the OLMo framework (allenai, 2024) to pretrain StackTrans. Our corpora comes from Dolma (Soldaini et al., 2024) and Smoll (Allal et al., 2025), which contain high-quality natural language, math, and Python code examples with diverse domains. We carry out data filtration, ultimately obtaining approximately 980 billion tokens. To scale up model parameters, we adapt the Dolma v1.6-sample configuration in OLMo, using roughly 80 billion tokens for each variant model training. For scaling up training tokens, we train StackTrans with 360M parameters on a sampled subset of 200 billion tokens from pretraining corpora.
@@ 6 Discussion
We evaluate all the variants introduced above on the V2 and V3 validation sets (Zhu et al., 2024). Experimental details are provided in §G. The ablation results are presented in Table 3. The QueueTrans variant exhibits notably higher perplexity and lower overall accuracy compared to StackTrans, particularly on tasks involving hierarchical or recursive patterns. This outcome is consistent with our expectation that queue operations are inherently less effective at modeling nested dependencies and grammars. Similarly, the push-only variant performs poorly, with both training and validation losses significantly deteriorating. The absence of pop operations impairs its ability to dynamically manage and retrieve stored information, thereby reducing its overall effectiveness. Both single-head and full-dimensional variants are consistently outperformed by StackTrans. The multi-head mechanism enhances flexibility by enabling parallel decomposition of stack streams, while the low-rank mechanism reduces computational costs without compromising much modeling capacity.
@@
In addition to the configuration of the standard Transformer layers, StackTrans has three key hyperparameters – the number of stack heads $H$, the dimension of each stack head $d_{s}$, and the stack size $S$. We perform grid search over reasonable ranges for these hyperparameters, training StackTrans with 20 billion tokens. The curves of training loss and validation loss (evaluated on V2 (Zhu et al., 2024)) are plotted in Figure LABEL:fig:stackablation. It is clear that $H$ is crucial for parallelism, but Figure LABEL:fig:ablation:head indicates that performance plateaus once $H$ surpasses a certain threshold. From Figure LABEL:fig:ablation:dim, we observe that setting $d_{s}$ within the range from 16 to 64 balances the computational cost and the model’s expressiveness effectively. Similarly, Figure LABEL:fig:ablation:slots shows that increasing $S$ from 24 to 32 has nearly no impact on performance. Given that a larger $S$ leads to higher storage overhead, we make a trade-off and ultimately set $S$ to 24 during our evaluation.